miracle711:

Ronnie, Mack, Kim, and Ada together stacked a total of 40 pizzas in a pan. Ronnie stacked 0.25 of the pizzas, Mack stacked 9 of the 40 pizzas, Kim stacked 35% of the pizzas, and Ada stacked the remaining pizzas. Who stacked the most pizzas? (5 points)

CourtneyWatson:

Which one do you think incorrect?

miracle711:

its kim

CourtneyWatson:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @miracle711 its kim \(\color{#0cbb34}{\text{End of Quote}}\) Yes you are correct

Katrin369:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @miracle711 its kim \(\color{#0cbb34}{\text{End of Quote}}\) yup you got it:)!

xxDeppressionxx:

incorrect

Katrin369:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx incorrect \(\color{#0cbb34}{\text{End of Quote}}\) what?

xxDeppressionxx:

Kim stacked the most

CourtneyWatson:

Yes

Katrin369:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx Kim stacked the most \(\color{#0cbb34}{\text{End of Quote}}\) thats what she/he said

xxDeppressionxx:

they asked which one was incorrect

xxDeppressionxx:

It said who stacked the most pizza's

xxDeppressionxx:

So the correct answer would have bee kim stacked the most pizza's

Katrin369:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx So the correct answer would have bee kim stacked the most pizza's \(\color{#0cbb34}{\text{End of Quote}}\) ya

jhonyy9:

Ronnie, Mack, Kim, and Ada together stacked a total of 40 pizzas in a pan. Ronnie stacked 0.25 of the pizzas, Mack stacked 9 of the 40 pizzas, Kim stacked 35% of the pizzas, and Ada stacked the remaining pizzas. Who stacked the most pizzas? total there are 40 pizzas Ronnie stacked 0,25 so this is 25/100 = 25% Mack stacked 9 of the 40 40 mean 100% 9 mean x x = 9*100/40 = 22,5% Kim stacked 35% Ada stacked the remaining part so 25% + 22,5% + 35% = 82,5% and 100 -82,5 = 17,5 so Ada stacked 17,5% and now from these just you need select the right answer hope helped understandably easy